OpenStudy (anonymous):

How do you convert the polar point (7, 5pi/4) to its cartesian form?

jimthompson5910 (jim_thompson5910):

by using these two formulas x = r*cos(theta) y = r*sin(theta)

jimthompson5910 (jim_thompson5910):

you are given r = 7 and theta = 5pi/4

jimthompson5910 (jim_thompson5910):

the cartesian or rectangular form of the point is (x,y)

OpenStudy (anonymous):

so x=-4.96 and y=-4.96?

OpenStudy (anonymous):

or \[-7(\sqrt{2})/2 , -7(\sqrt{2})/2 \]

jimthompson5910 (jim_thompson5910):

go with the exact form unless they ask for the approximate

jimthompson5910 (jim_thompson5910):

but yeah that's correct

OpenStudy (anonymous):

Ok, thanks again

jimthompson5910 (jim_thompson5910):

np